### v1

# HEST-1k: A Dataset for Spatial Transcriptomics and Histology Image Analysis

Abstract
Spatial transcriptomics (ST) enables interrogating the molecular composition of tissue with ever-increasing resolution, depth, and sensitivity. However, costs, rapidly evolving technology, and lack of standards have constrained computational methods in ST to narrow tasks and small cohorts. In addition, the underlying tissue morphology as reflected by H&E-stained whole slide images (WSIs) encodes rich information often overlooked in ST studies. Here, we introduce HEST-1k, a collection of 1,108 spatial transcriptomic profiles, each linked to a WSI and metadata. HEST-1k was assembled using HEST-Library from 131 public and internal cohorts encompassing 25 organs, two species (Homo Sapiens and Mus Musculus), and 320 cancer samples from 25 cancer types. HEST-1k processing enabled the identification of 1.5 million expression–morphology pairs and 60 million nuclei. HEST-1k is tested on three use cases: (1) benchmarking foundation models for histopathology (HEST-Benchmark), (2) biomarker identification, and (3) multimodal representation learning. HEST-1k, HEST-Library, and HEST-Benchmark can be freely accessed via https://github.com/mahmoodlab/hest.
**footnotetext: Equal contribution
1 Introduction
Advances in molecular profiling enable spatially-resolved gene expression analysis with increasingly large gene panels, enhanced spatial resolution, and greater sensitivity Asp et al. (2020); Rao et al. (2021). From the early days of bulk RNA sequencing constrained by its coarse resolution and limited gene panels, spatially-resolved technologies have progressed to achieve whole-transcriptome sequencing at sub-cellular resolution Li and Wang (2021). In cancer research, spatial transcriptomics (ST) holds particular promise for characterizing the tumor microenvironment, a key element in understanding disease progression and treatment response Ståhl et al. (2016b); Marx (2021); de Visser and Joyce (2023); Xiao and Yu (2021). With the large amount of transcriptomics data generated by a single ST sample (e.g., $>$10 million transcripts are detected in a typical 10x Genomics Xenium assay), computational methods are often used to identify promising targets and biomarkers, such as based on gene expression clustering or gene-set enrichment analysis Rao et al. (2021).
However, high costs and rapidly evolving technology have constrained computational methods to narrow tasks and data cohorts of only few patients Meylan et al. (2022); Janesick et al. (2023); Valdeolivas et al. (2023). Consequently, we observe a lack of standardized resources and unified formats for storing ST, which limits the development of deep learning models on a large scale Moses and Pachter (2022). In addition, the underlying tissue morphology, traditionally visualized in hematoxylin and eosin (H&E)-stained tissue sections (whole-slide images, WSIs), is often overlooked in ST studies, despite encoding valuable information. In particular, pairs of ST and WSI enable analyzing expression changes in their morphological context, which may facilitate the discovery of morphological biomarkers (e.g., changes in nuclear shape) that correspond to gene regulation patterns. In translational medicine, ST could be employed for identifying biomarkers, which are then morphologically validated and characterized by widely deployed clinical modalities such as H&E. Alternatively, pairs of ST and WSI can enable multimodal tissue representation learning for joint modeling of the morphomolecular signature of tissue at a scale and resolution beyond bulk RNA sequencing Chen et al. (2022). Finally, the development of “foundation models” for encoding histopathology image patches Wang et al. (2021); Filiot et al. (2023); Chen et al. (2024); Huang et al. (2023) has increased the need for new, diverse, and challenging benchmarks beyond diagnostic tasks. Using ST, new tasks can be defined for predicting gene expression from histology.
Here, we introduce HEST-1k, a collection of paired ST and H&E-stained WSIs curated from public and internal cohorts (Figure 1.a). HEST-1k comprises 1,108 samples from 131 cohorts encompassing 25 organs, two species (Homo Sapiens and Mus Musculus), and 320 cancer samples from 25 subtypes. Processing all samples in HEST-1k resulted in 1.5 million expression–morphology pairs and 60 million detected nuclei. With new cohorts frequently made public, we also introduce the HEST-Library, a Python package for querying HEST-1k and assembling new legacy data (Figure 1.b). We highlight the potential of HEST-1k through three use cases: (i) benchmarking foundation models for histology using the HEST-Benchmark, a set of ten tasks (nine human cancer types and ten organs) for gene expression prediction from histology and evaluated on ten state-of-the-art models (Figure 1.c), (ii) a proof-of-concept demonstrating the use of HEST-1k for biomarker characterization (Figure 1.d), and (iii) a proof-of-concept for expression-guided fine-tuning of foundation models for histology (Figure 1.e).
Figure 1: Overview of HEST. a. Description of HEST-1k, a dataset of $n$=1,108 paired Spatial Transcriptomics, H&E-stained whole-slide images and metadata. b. Overview of HEST-Library functionalities. c., d., e. Potential applications of HEST-1k include benchmarking foundation models for histology (c.), biomarker discovery (d.) and multimodal representation learning (e.).
2 Related work
Libraries for ST analysis. Libraries to process, visualize, and analyze ST have been built around two core pipelines: Scanpy Wolf et al. (2018) (and the Anndata format) in Python and Seurat Butler et al. (2018) in R. Scanpy has served as the foundation for several subsequent developments such as Squidpy Palla et al. (2022) for spatial data exploration at cellular-, gene-, and morphological-level, SpatialData Marconato et al. (2024) for multi-technology integration and deep learning interfacing, and STlearn Pham et al. (2023) for cell-cell interactions and spatiotemporal trajectory analyses. In R, Seurat Butler et al. (2018) has been consolidated with packages such as BayeSpace Zhao et al. (2021) for clustering and spot super-resolution, and Giotto suites Chen et al. (2023) for preprocessing, data integration, and visualization of multiple ST technologies. 10x genomics also includes software analytics through the Xenium and Visium Explorer pipelines for multimodal visualization, nuclear segmentation, and cell deconvolution. However, none of these pipelines were designed to handle the diversity of legacy data, where datasets can suffer from missing or incorrect data, such as alignment mismatches, incorrect pixel resolution, inconsistent image file formats, etc.
Molecular profile prediction from H&E. Molecular profiling from histology images has been explored both at (1) slide-level to predict bulk molecular status/changes from a WSI, and at (2) patch-level to predict local molecular status/changes from regions-of-interest. (1) Slide-level profiling has been proposed to predict the presence of gene mutations Saldanha et al. (2023); Wagner et al. (2023); Kather et al. (2020); Echle et al. (2021); Fu et al. (2020); Wang et al. (2019); Loeffler et al. (2022), microsatellite instability Wagner et al. (2023); Kather et al. (2019b), and gene expression changes Schmauch et al. (2020); El Nahhas et al. (2024); He et al. (2020); Alsaafin et al. (2023), among others. The motivation is two-fold: First, patient screening to substitute or emulate costly clinical molecular assays, and second, to identify morphological correlates of molecular alterations for discovering novel biomarkers. Such studies can be conducted on large patient cohorts as they mainly rely on data generated by the routine clinical workflow (e.g., using TCGA cohorts with $>$11,000 cases from 33 cancer types). (2) With ST, several works have explored predicting expression changes from regions-of-interest He et al. (2020); Xie et al. (2023); Pang et al. (2021); Zhao et al. (2024); Rahaman et al. (2023); Monjo et al. (2022); Mondol et al. (2023); Chung et al. (2024); Wang et al. (2024). Due to limited cohort sizes (typically one to ten patients), transfer learning has become the norm using pretrained networks based on ConvNets He et al. (2020) or Vision Transformers Xie et al. (2023); Pang et al. (2021). Transcriptomic measurements are inherently noisy, hence the development of context-based methods that integrate global and local information from surrounding ST spots Chung et al. (2024); Zhao et al. (2024); Wang et al. (2024). While recent technologies offer near-single-cell resolution (such as Visium HD and Xenium), legacy assays operate at a more coarse resolution, which can be upsampled using super-resolution techniques Zhao et al. (2021); Bergenstråhle et al. (2022); Zhang et al. (2024). The potential clinical and research implications of such methods are still being explored, with HEST-1k potentially catalyzing their large-scale development.
Foundation models in pathology. A fundamental task in computational pathology is to extract general-purpose embeddings of image patches (typically 256$\times$256-pixel regions) that can then be used for downstream tasks, such as diagnosis or prognosis prediction. To achieve this, self-supervised learning (SSL) has been extensively applied  Koohbanani et al. (2021); Filiot et al. (2023); Vorontsov et al. (2023); Wang et al. (2021); Kang et al. (2023); Ciga et al. (2022); Wang et al. (2022); Azizi et al. (2023); Chen et al. (2024); Lu et al. (2024), such as based on the DINOv2 framework Oquab et al. (2023). Models for general-purpose patch encoding are trained on increasingly large and more diverse patient cohorts with models growing in model size (e.g., Virchow uses a ViT-Huge trained on more than 1.5M WSIs Vorontsov et al. (2023)). As the number of such models rapidly increases, new, diverse, and challenging benchmarks are needed to replace or complement well-established tasks where performance has saturated. HEST-Benchmark aims to address this by offering a set of ten patch-level tasks for gene expression prediction from histology.
Patch-level benchmarks in histopathology. Early task and dataset contributions in computational pathology revolved around classifying small regions of interest. Over the years, a variety of benchmarks have been established: In prostate cancer, Gleason grading at pixel- and patch-level has been widely explored, with public resources such as PANDA Bulten et al. (2022), AGCC Huo et al. (2022), DiagSet Koziarski et al. (2024), and SICAPv2 Silva-Rodríguez et al. (2020). In colorectal cancer, datasets have been proposed for tissue classification, such as HunCRC Pataki et al. (2022), UniToPatho Barbano et al. (2021), MHIST Wei et al. (2021), and CRC-100k Kather et al. (2019a). In breast cancer, morphological subtyping has been studied (e.g., for atypical ductal hyperplasia detection), such as BACH Aresta et al. (2019), BRACS Brancati et al. (2021), and BreakHis Spanhol et al. (2015), and for lymph node metastasis detection with Patch CAMELYON (pCAM) Veeling et al. (2018), respectively. However, the performance on many of these datasets has saturated; for instance, Gleason scoring reaches similar or better performance than pathologists Bulten et al. (2022), which limits objective comparisons of new methods and hinders well-informed model selection for developing better features. Instead, HEST-Benchmark provides a collection of diverse and challenging tasks that enable assessing the predictive capabilities of foundation models for histology.
3 HEST-1k Dataset
We present HEST-1k, a dataset of paired ST, H&E-stained WSIs, and metadata (Figure 1.a). To this end, we extracted all publicly available cohorts that provide ST with H&E-stained whole-slide images. Specifically, we harvested data from 10x Genomics public datasets (TENX) [^1] https://www.10xgenomics.com/datasets, Mendeley (MEND) [^2] https://data.mendeley.com/, Spatial-Research (SPA) [^3] https://www.spatialresearch.org/, Zenodo (ZEN) [^4] https://zenodo.org/, the National Center for Biotechnology Information (NCBI) [^5] https://www.ncbi.nlm.nih.gov/gds/, GitHub [^6] https://github.com/, the Human Cell Atlas [^7] https://data.humancellatlas.org/, and internal data cohorts. A summary of all sources is provided in Appendix Table A1 with specifics in Appendix Table A2, A4, A6,A7,A8,A9, and A10.
3.1 Metadata
As Spatial Transcriptomics were not intended for large-scale computational research, they are provided in various formats (e.g., images can be in JPG or TIFF format, with or without cross-modal alignment files) and resolutions. We unified all data with comprehensive metadata with generic-, histology-, and expression-related descriptors for all samples. Generic: We provide the reference to the original publication, download link, year of publication, license, and sample species. Each sample is then categorized as either healthy, cancer, treated (which refers to a post-compound administration), genetically modified (mostly knock-out mouse samples), or other diseases (with an additional description). All cancer samples were unified using the OncoTree code, a taxonomy of cancer types provided by the Memorial Sloan Kettering Cancer Center [^8] https://oncotree.mskcc.org. Finally, we include the organ using the highest level of the OncoTree taxonomy as a reference. Expression: We include the number of genes and spots per sample, the spot resolution and spacing, the total number of reads, and the mean number of reads per spot. We additionally provide the transcriptomic technology (ST, Visum, Visium HD, or Xenium). Histology: We provide the image resolution (in $\mu$m/pixel) and magnification as 10$\times$ (1.15 to 0.8 $\mu$m/px), 20$\times$ (0.8 to 0.4 $\mu$m/px) and 40$\times$ (0.4 to 0.1 $\mu$m/px). All images with a resolution lower than 1.15 $\mu$m/px were discarded to ensure an acceptable image quality. In addition, we provide the image size at the highest resolution, and tissue preparation protocol (frozen or formalin-fixed paraffin-embedded, FFPE).
3.2 Histology
All tissue sections were normalized and transformed into a generic TIFF object, a pyramidal image that can easily be integrated into computational frameworks using Openslide or viewers such as QuPath Bankhead et al. (2017). In addition, we provide a contour object that delineates all the tissue regions identified in the image. We developed a robust tissue detection method where we fine-tuned a DeepLabV3 Chen et al. (2017) model with an ImageNet-pretrained ResNet50 backbone on a set of annotated segmentation regions. From the tissue segmentation, we extracted 224$\times$224-pixel patches at 20$\times$ magnification around each spot. This yielded 1.5 million valid patches for which a corresponding expression profile was derived. Such patching can readily be used for various downstream tasks, such as employed in the HEST-Benchmark or multimodal fine-tuning of foundation models for histology (Section 5 and 7).
3.3 Nuclear segmentation and classification
In addition to patching, we include nuclear segmentation that delineates each nucleus identified in all slides from HEST-1k. Specifically, we use CellViT Hörst et al. (2024), a state-of-the-art nuclear segmentation model that was trained on the PanNuke dataset Gamper et al. (2019, 2020). CellViT enables joint instance segmentation and classification of each nucleus into five classes: neoplastic epithelial, non-neoplastic epithelial, inflammatory, stromal, and necrotic. On average, we identified 55.5k nuclei per slide (standard deviation 91.3k nuclei), for a total of 59.6 million nuclei identified across all samples. Among those 17.6 million are classified as neoplastic, 21.5 million as stromal, 4.9 million as normal epithelial, 15.4 million as inflammatory, and 76 thousand as necrotic. The resulting nuclear segmentation and classification can easily be visualized using QuPath (using geojson) or loaded as Python/R objects (using json).
3.4 Gene expression
All expression data were unified in a Anndata object that can be loaded with scanpy. Anndata encodes the gene names (as var) and number of spots (as obs). Each entry represents the raw transcript counts of a gene in a given spot. No additional normalization was conducted, and we let users explore various normalization strategies based on needs, e.g., using total count normalization, log-based re-weighting, etc. In addition, we include metadata to specify the number of genes, the gene panel, and the tissue site.
To use the expression in tandem with the WSI, an alignment file describing the mapping between the image and the spots is needed. However, relying on publicly available alignment information brings three challenges: (1) most datasets report alignment with respect to a low-resolution version of the image, (2) they are not standardized, and (3) alignment quality can be low. To address these limitations, we re-aligned all samples under the same unified format between the WSI and the corresponding expression profile. To this end, we developed an automatic alignment pipeline (see Section 4) and embedded the alignment in the scanpy object.
4 HEST-Library
The HEST-Library is built around scanpy and Anndata. At its core, the HEST-Library enables (1) assembling and querying HEST-1k, and (2) running the HEST-Benchmark (see Section 5). We describe its core functionalities, in particular for unifying legacy data.
Conversion to generic TIFF. We integrate functions to convert a WSI from common formats found in public ST datasets (e.g., OME.TIF, JPG, BigTIFF, etc.) to a pyramidal generic TIFF format. Pyramidal formats offer seamless integration with OpenSlide (commonly used in computational pathology pipelines) and QuPath (open-access software for WSI visualization and annotation).
Automatic alignment. Spot alignment is crucial to ensure an accurate match between the ST spots and the WSI. While software such as LoupeBrowser enables manual alignment using the fiducials (in Visium), the process remains time-consuming when processing large batches of samples. Instead, we implemented an automatic fiducial detection algorithm based on Yolov8 Redmon et al. (2016) for processing Visium samples (Appendix Figure 4). Specifically, we manually annotated 119 fiducial regions that we further augmented using tissue and fiducials mixing. We then fine-tuned Yolov8 pretrained on COCO dataset. In samples without fiducials (e.g., early versions of Spatial Transcriptomics), we realigned using the provided spot position files.
Automatic detection of image resolution. From the alignment and the spot resolution, we can infer the exact pixel resolution. We start by deriving the distance in pixel between two neighboring spots and leverage the known inter-spot distance in $\mu$m to estimate the pixel width in $\mu$m/px. For Xenium samples, we directly read the H&E alignment file which provides an affine transformation from the DAPI-stained image (with known pixel size) to the H&E image. We then compared the self-reported image resolution and our re-estimations to manually inspect and correct discrepancies.
Conversion to Anndata. ST data is provided in multiple formats, such as CSVs, MEX, TXT, h5, etc. We provide functions to unify a large set of existing formats into a AnnData object that stores the raw transcript counts as a matrix of genes by the number of spots.
Tissue segmentation and patching. We provide a tissue segmentation pipeline optimized for Visium/Xenium images. The segmented tissue can automatically be tessellated into fixed-size image patches at a predefined resolution around each spot.
Automatic HEST-1k download. To facilitate downloading part or all of the HEST-1k dataset (in total 825 GB), we implemented an easy download option where the user can specify entries of the metadata, for instance, to query all human breast cancer cases.
5 HEST-Benchmark
From HEST-1k, we curated the HEST-Benchmark, a set of ten tasks for gene expression prediction from histology in human cancer samples. The goal is two-fold: (i) benchmarking foundation models for histology under a diverse and challenging benchmark, and (ii) understanding the predictive capabilities of state-of-the-art models in predicting expression from morphology. Compared to existing tasks (e.g., Camelyon16), HEST-Benchmark brings more morphological and data acquisition diversity, and more complex tasks as expression prediction is inherently challenging.
5.1 Task definition
We define ten tasks with data from nine human cancers and ten organs (nine primary and one metastatic dataset), which include invasive ductal carcinoma (breast cancer, IDC, Task 1), prostate adenocarcinoma (prostate cancer, PRAD, Task 2), pancreatic adenocarcinoma (pancreatic cancer, PRAD, Task 3), skin cutaneous melanoma (skin cancer, SKCM, Task 4), colonic adenocarcinoma (colon cancer, COAD, Task 5), rectal adenocarcinoma (rectum cancer, READ, Task 6), clear cell renal cell carcinoma (kidney cancer, ccRCC, Task 7), hepatocellular carcinoma (liver cancer, HCC, Task 8), lung adenocarcinoma (lung cancer, LUAD, Task 9), and axillary lymph nodes in IDC (metastatic, LYM-IDC, Task 10). Additional information is provided in Appendix Table A11.
For each task, we predict the expression of the top 50 genes with the highest normalized variance across all samples from 112$\times$112 $\mu$m H&E regions (equivalent to 224$\times$224-pixel patches at 20$\times$). To avoid train/test patient-level data leakage, we use patient-stratified splits, resulting in a $k$-fold cross-validation, where $k$ is the number of patients. In ccRCC, we use $k/2$-fold cross-validation due to the large number of patients.
5.2 Evaluating foundation models for histology on HEST-Benchmark
We use the HEST-Benchmark to evaluate ten patch encoders for histology. Namely, ResNet50 (IN) Lu et al. (2020) (ImageNet pretrained), KimiaNet Riasatian et al. (2021) (pretrained with supervised subtyping classification), Ciga Ciga et al. (2022) (SimCLR pretrained on public histology data), CTransPath Wang et al. (2022) (adapted MoCov3 pretrained on TCGA and PAIP), Remedis Azizi et al. (2023) (iBOT pretrained on TCGA), Phikon Filiot et al. (2023) (iBOT pretrained on TCGA), PLIP Huang et al. (2023) (visual-language model based on CLIP pretrained on histology data from Twitter), UNI Chen et al. (2024) (DINOv2 ViT-Large pretrained on public and private histology data), CONCH Lu et al. (2024) (visual-language model using CoCa pretrained on captions from publications and educational resources), and GigaPath Xu et al. (2024) (DINOv2 ViT-Giant pretrained on proprietary data). Additional information is provided in Table A12 and Section C.3.
Table 1: HEST-Benchmark results. Random Forest regression evaluation (70 trees) on publicly available state-of-the-art patch encoders. Model performance measured with Pearson correlation. We report the mean $\pm$ standard deviation computed over all folds (or patients). Best is bold, second best is underlined. Models sorted by publication date.
ResNet50 KimiaNet Ciga CTransPath Remedis Phikon PLIP UNI CONCH GigaPath
IDC 0.440 0.420 0.406 0.454 0.491 0.430 0.436 $\underline{0.502}$ $\bm{0.504}$ 0.492
±$\pm 0.03$$\pm 0.04$$\pm 0.02$$\pm 0.04$$\pm 0.06$$\pm 0.06$$\pm 0.04$$\pm 0.06$$\pm 0.05$$\pm 0.05$
PRAD 0.318 0.328 0.332 0.346 0.335 $\bm{0.377}$ 0.362 0.357 $\underline{0.373}$ 0.372
±$\pm 0.02$$\pm 0.01$$\pm 0.00$$\pm 0.01$$\pm 0.00$$\pm 0.00$$\pm 0.02$$\pm 0.00$$\pm 0.03$$\pm 0.01$
PAAD 0.389 0.410 0.397 0.406 $\bm{0.451}$ 0.372 0.392 0.424 $\underline{0.431}$ 0.425
±$\pm 0.07$$\pm 0.08$$\pm 0.07$$\pm 0.08$$\pm 0.06$$\pm 0.06$$\pm 0.09$$\pm 0.06$$\pm 0.05$$\pm 0.07$
SKCM 0.446 0.452 0.484 0.535 0.577 0.516 0.461 $\bm{0.613}$ $\underline{0.582}$ 0.541
±$\pm 0.06$$\pm 0.03$$\pm 0.01$$\pm 0.06$$\pm 0.00$$\pm 0.04$$\pm 0.06$$\pm 0.02$$\pm 0.04$$\pm 0.03$
COAD 0.107 0.080 0.102 0.123 0.125 0.137 0.112 $\bm{0.147}$ 0.124 $\underline{0.139}$
±$\pm 0.06$$\pm 0.03$$\pm 0.04$$\pm 0.05$$\pm 0.06$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.03$
READ 0.051 0.038 0.046 0.083 0.099 0.138 0.063 $\bm{0.162}$ 0.132 $\underline{0.156}$
±$\pm 0.08$$\pm 0.05$$\pm 0.09$$\pm 0.09$$\pm 0.11$$\pm 0.10$$\pm 0.10$$\pm 0.08$$\pm 0.08$$\pm 0.08$
CCRCC 0.136 0.136 0.127 0.171 $\bm{0.200}$ 0.178 0.124 $\underline{0.186}$ 0.149 0.182
±$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.06$$\pm 0.05$
HCC 0.034 0.028 0.045 $\bm{0.060}$ $\underline{0.059}$ 0.041 0.038 0.051 0.040 0.055
±$\pm 0.01$$\pm 0.01$$\pm 0.00$$\pm 0.00$$\pm 0.00$$\pm 0.02$$\pm 0.00$$\pm 0.00$$\pm 0.01$$\pm 0.01$
LUNG 0.497 0.507 0.515 0.531 $\bm{0.573}$ 0.541 0.533 0.511 $\underline{0.569}$ 0.547
±$\pm 0.01$$\pm 0.01$$\pm 0.02$$\pm 0.02$$\pm 0.02$$\pm 0.03$$\pm 0.02$$\pm 0.03$$\pm 0.02$$\pm 0.03$
LYMPH_IDC 0.205 0.206 0.218 0.238 0.243 0.243 0.229 0.234 $\bm{0.249}$ $\underline{0.248}$
±$\pm 0.05$$\pm 0.06$$\pm 0.07$$\pm 0.06$$\pm 0.06$$\pm 0.06$$\pm 0.06$$\pm 0.05$$\pm 0.06$$\pm 0.05$
Average 0.262 0.261 0.267 0.295 0.315 0.297 0.275 $\bm{0.319}$ 0.315 $\underline{0.316}$
We learn a regression model to map model-specific patch embeddings (512 to 2048 dimensions) to the log1p-normalized expression of the top-50 highly variable genes. All tasks are evaluated using the Pearson correlation between the predicted and measured gene expression. We report mean and standard deviation across all folds (or patients). All experiments were run on a single NVIDIA 3090 GPU. Table 1 summarizes HEST-Benchmark when employing a Random Forest regression model (70 trees), and Table A13 when employing a Ridge regression model.
When using Random Forest, UNI Chen et al. (2024), GigaPath Xu et al. (2024), Remedis Azizi et al. (2023), and CONCH Lu et al. (2024) provide similar performance (0.315 to 0.319 average Pearson correlation across all tasks), outperforming all other baselines. When using a Ridge regression, CONCH Lu et al. (2024) significantly outperforms all models with an average of +5.0% absolute performance gain over the second-best model, Phikon, and 5.3% better than the third-best model, CTransPath with an average Pearson correlation of 0.316.
Supervised vs. contrastive vs. student-teacher pretraining: When employing both Random Forest or Ridge regression, supervised models (ResNet50-IN and KimiaNet) exhibit lower performance than other models. This can be attributed to (1) a lack of sufficiently large and diverse pretraining datasets and models or (2) fundamental limitations of supervised models in transferring to challenging tasks. Using Random Forest, the best student-teacher models (UNI, CONCH, GigaPath) provide similar performance to the best contrastive approach (Remedis).
Vision Transformer Base vs. Large vs. Giant: CONCH (ViT-Base, 86M parameters), UNI (ViT-Large, 307M) and GigaPath (ViT-Giant, 1.13B) provide similar performance despite GigaPath using 13$\times$ more parameters than CONCH. UNI and GigaPath have been trained in a unimodal fashion by building “views” using traditional augmentation techniques, such as color jittering and cropping. In contrast, CONCH was trained in two steps: first using unimodal iBOT pretraining, and then by fine-tuning the resulting model on image-caption pairs. Such fine-tuning step could explain its parameter-efficiency and high performance in HEST-Benchmark.
Random Forest vs. Ridge regression: The best-performing model using Random Forest (UNI, 0.319) is similar to the best-performing model using Ridge regression (CONCH, 0.316), showing the predictive capabilities of both regression models. However, Ridge regression leads to poor performance using larger models (GigaPath, Remedis, and UNI), which could be explained by the curse of dimensionality (CONCH embedding dimension is 3$\times$ smaller than GigaPath). We also emphasize that the performance could be optimized using more complex models (such as MLP) or dedicated hyper-parameter search.
Overall, HEST-Benchmark brings new insights into the performance of patch encoders for histology. We observe that (1) Random Forest regression leads to multiple models performing similarly (CONCH, UNI, GigaPath) which may suggest a form of plateauing of patch encoders, (2) CONCH is the only model that shows high performance using both Ridge and Random Forest regression models, which could be attributed to its compact size (ViT-B) and visual-language fine-tuning step, and (3) performance remains low for some tasks (e.g., READ and HCC) which implies that for certain cancer types, the morphology might not be as reflective of gene expression, as well as leaving room for follow-up works.
6 HEST for discovery
Figure 2: HEST for biomarker discovery: Analysis of an invasive ductal carcinoma Xenium sample. a. IDC Xenium samples with neoplastic nuclei overlaid in red ($n_{c}=168,033$ detected nuclei). Gray scale bar represents 2 mm. b. Heatmap of Xenium expression of gene GATA3. Blue and red values indicate above and below the mean (in white), respectively. c. Heatmap of neoplastic nuclear area. d. Four randomly selected regions with CellViT segmentation of the neoplastic nuclei. Black scale bar represents 30 $\mu$m. e., f. Correlation between area and GATA3, and minor axis length and MYBPC1.
HEST-1k also enables the analysis of interactions and correlations between tissue morphology (as seen in H&E) and local gene expression (as provided in ST). Here, we showcase the capabilities of HEST-1k (1) by studying morphological correlates of expression changes in invasive breast cancer and (2) by visualizing tumor heterogeneity both on the morphological and molecular sides. Specifically, we focus on invasive ductal carcinoma (IDC) samples imaged with Xenium. Using CellViT nuclear segmentation and classification, we identified neoplastic nuclei (exemplified in two samples: Figure 2.a with n=168,033 nuclei and Appendix Figure 5.a with n=342,018 nuclei). We then overlay the WSI with the expression of specific genes, such as GATA3, a known prognostic gene in breast cancer Mehra et al. (2005)(Figure 2.b). This qualitatively shows that high GATA3 expression is associated with cancerous regions and reveals heterogeneity within invasive regions (e.g., the right-most region shows higher expression of GATA3 than the rest of the tumor, Figure 2.b). Using the nuclear segmentation, we can compute human-interpretable features related to nuclear size (area, perimeter, major axis length, minor axis length, and equivalent diameter), topology and shape (roundness, ellipticity, eccentricity, extent, and roughness), and cell distribution (cell density and crowdedness). A heatmap of the nuclear area of neoplastic cells also indicates morphological heterogeneity among neoplastic regions (Figure 2.c,d). Regions with a high nuclear area and elevated GATA3 expression notably overlap, suggesting that this tumor exhibits molecular heterogeneity, which to some degree is morphologically expressed.
To investigate this hypothesis, we measured the Pearson correlation between the expression of GATA3 and nuclear area in neoplastic cells (Figure 2.e). We observe a significant correlation (R=0.47), which is also observed in other genes and morphological features (Figure 2.f). Overall, out of the 12 human-interpretable features we analyzed, we found the highest association with gene expression for size-related features, while features involving topology, shape, and cell distribution had a lower correlation (R<0.2). A similar analysis in another IDC sample (Appendix Figure 5.b,c) confirmed these observations. In particular, we found the highest associations between nuclear size and expression for the genes FLNB (R=0.49) and TPD52 (R=0.49), both involved in breast tumor growth and proliferation Bandaru et al. (2014); Ren et al. (2021), and FOXA1 (R=0.47), a known prognostic factor associated with better survival Badve et al. (2007); Wolf et al. (2007).
Such analysis highlights how HEST-1k can be used for identifying fine-grained morphological correlates of expression. Additional analyses of morphological and molecular tumor heterogeneity can be leveraged at larger scale using similar approaches.
7 HEST for multimodal representation learning
Access to spatially-resolved expression–morphology pairs unlocks new directions for multimodal representation learning. Several problem statements can be explored, such as cross-modal alignment and retrieval, multimodal fusion, etc. Here, we fine-tune CONCH (strong performer on HEST-Benchmark) on five Xenium invasive breast cancer cases (four ductal and one lobular case) using contrastive alignment. We hypothesize that the resulting breast cancer-specific patch encoder, termed CONCH-FT, can better encode the underlying molecular landscape associated with disease-specific morphologies. To validate the hypothesis, CONCH-FT is benchmarked on an independent breast cancer cohort for molecular subtyping against its non-finetuned version.
[…]
We evaluate the resulting CONCH-FT model to predict ER, PR, and HER2 expression status (binary) from WSIs in the BCNB dataset Xu et al. (2021) (n=1,058 WSIs). To generate a slide representation for a WSI, we take the average of the patch embedding in the WSI, which is subsequently mapped to the expression status using logistic regression (Table 2). We observe that CONCH-FT outperforms CONCH on most metrics, demonstrating that pan-tissue histology patch encoders can be further fine-tuned to obtain better tissue-specific patch encoders. This is further validated by the larger rank induced by the patch embedding space Garrido et al. (2023) for CONCH-FT, suggesting better expressivity of the patch embeddings. While these results are based on only five WSIs, we anticipate additional benefits when training with larger disease-specific cohorts.
Table 2: CONCH fine-tuning on invasive breast cancer. Logistic regression evaluation for ER/PR/HER2 expression status on BCNB (n=1,058 WSIs). A WSI is represented by the average of the patch embeddings within each WSI. We report the mean ± standard deviation computed over all folds (or patients) for ROC-AUC (AUC) and balanced accuracy (Bal.acc.). Best is bold.
Rank ER PR HER2
AUC Bal.acc. AUC Bal.acc. AUC Bal.acc.
CONCH 144.66 0.881 0.745 0.810 0.698 0.715 $\bm{0.624}$
CONCH-FT $\bm{146.47}$ $\bm{0.884}$ $\bm{0.752}$ $\bm{0.818}$ $\bm{0.714}$ $\bm{0.724}$ 0.615
8 Discussion
Summary. We assembled HEST-1k, a dataset comprising paired spatial transcriptomics, H&E-stained whole-slide images, and comprehensive metadata built from public and internal cohorts. HEST-1k includes 1,108 samples, encompassing 1.5 million spots and over 60 million cells. The scale and comprehensiveness of HEST-1k, supported by the HEST-Library, enable exploring directions such as biomarker discovery and multimodal representation learning. Additionally, motivated by the need for new diverse and challenging patch-level benchmarks, we curated the HEST-Benchmark, a set of ten tasks covering nine cancer types and ten organs for gene expression prediction from histology. HEST-Benchmark revealed new insights into the predictive capabilities of foundation models for histology, such as the added benefits of vision-language alignment Lu et al. (2024).
Limitations. Our study includes limitations. First, research data, such as those generated in spatial transcriptomic, are inherently noisy. While we tried to minimize “label” noise (e.g., by re-estimating image magnification and alignment, and unifying cancer samples using oncotree code taxonomy), staining and compression artifacts, varying acquisition protocols, among others can negatively impact the quality of HEST-1k. Second, batch effects (on both the imaging and transcriptomic sides) can be significant across samples, datasets, and technologies. This study does not quantify these batch effects or explore methods to mitigate them. Lastly, although the HEST-Library was designed for versatility, it cannot cover all existing formats and should rather be viewed as a blueprint for processing ST data in a consistent and unified manner.
Future work. Spatial transcriptomics is rapidly evolving with new datasets frequently published. As they become available, we will keep updating HEST-1k with new resources. This study merely starts to uncover the potential of HEST-1k for advancing translational research and biomarker discovery, and we plan to explore these capabilities further. Additionally, the prospects for multimodal representation learning with HEST-1k are promising and are expected to grow with the addition of more data.
Acknowledgements
We thank Dr. Maxime Meylan for his insights and guidance on accessing data published in Meylan et al. (2022). S.J.W. was supported by the Helmholtz Association under the joint research school “Munich School for Data Science - MUDS” and the Add-on Fellowship of the Joachim Herz Foundation.
Checklist
  1. 1.
    Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] Each claim: HEST-1k, HEST-Library, HEST-Benchmark, HEST for discovery and multimodal fine-tuning are supported by dedicated sections in the main text, in addition to supplementary information provided in the appendix. In addition, all the code to reproduce results is made available.
  2. 2.
    Did you describe the limitations of your work? [Yes] We discuss limitations in the Discussion.
  3. 3.
    Did you discuss any potential negative societal impacts of your work? [Yes] We discuss potential negative societal impacts in the section Ethical considerations, intended usage, and license.
  4. 4.
    Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  5. 5.
    Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] In the abstract, we provide a link to access the HEST website and the HEST-Library on GitHub. HEST-Library includes a link to download all data to run the HEST-Benchmark. Finally, we provide all metadata associated with HEST-1k in a CSV as part of the supplementary material. Processed ST and WSIs will be made available upon acceptance.
  6. 6.
    Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] When relevant, we provide training details, such as in the HEST-Benchmark.
  7. 7.
    Did you report error bars? [Yes] HEST-Benchmark results include standard deviation computed from cross-validation across all patients.
  8. 8.
    Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
  9. 9.
    If your work uses existing assets, did you cite the creators? [Yes] All public resources used in this study are cited in Appendix Table A2, A4, A6, A7, A8, A9 and A10.
  10. 10.
    Did you mention the license of the assets? [Yes] Metadata associated with HEST-1k includes the license under which data were originally published. We ensured that the reported license allowed for distributing and creating derivatives of the data.
  11. 11.
    Did you include any new assets either in the supplemental material or as a URL? [Yes] As part of HEST-1k, we include internal datasets (see Appendix Table A9).
  12. 12.
    Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [No] We used public resources for which the license was allowing redistributing the work. Users are welcome to inspect the individual IRBs of each publicly available resource.
  13. 13.
    Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We manually ensured that none of the published and distributed data includes personally identifiable information or offensive content, such as personal health information.
Appendix A Ethical considerations, intended usage and license
All resources provided as part of this study are strictly for research purposes and must not be utilized to support any diagnostic procedures. Users are hereby notified that the nuclear segmentation and classification components are derived from a publicly available model. Consequently, this model should not be regarded as the definitive standard, and users should exercise particular caution when utilizing this part of the dataset. Despite our efforts to exclude sensitive information, such as patient names, addresses, and social security numbers, users are expressly prohibited from attempting to reverse engineer the data to extract any confidential patient information. On the presumption that users will adhere to the aforementioned restrictions, we have not identified any potential adverse societal impacts that could arise from using HEST-1k.
The dataset is hosted on HuggingFace Dataset webpage. All instructions are provided in the main README of HEST-Library. From there, users can choose to download HEST-1k in its entirety, or a subset (e.g., only breast cancer samples). HEST-1k, HEST-Benchmark, and HEST library are released under the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 Deed) [^9] https://creativecommons.org/licenses/by-nc-sa/4.0/.
Appendix B Background
This study connects two fields: (1) computational pathology, which primarily uses routinely acquired clinical data to determine outcomes such as disease diagnosis from H&E-stained digitized tissue sections, and (2) spatial transcriptomics, which has, so far, been confined to biological research and aims to identify new biomarkers predictive of disease progression or treatment response, among others.
B.1 Computational pathology
[…]
Such frameworks have been shown to achieve better or similar performance than humans for Gleason grading in prostate cancerBulten et al. (2022), metastasis detection in lymph nodesBejnordi et al. (2017), determining the origin of a cancer of unknown primaryLu et al. (2021a), predicting heart transplant rejectionLipkova et al. (2022), among others.
B.2 Spatial Transcriptomics (ST)
ST enables the measurement of gene activity and the mapping to its corresponding location on the tissue. In this study, we gather samples from two ST paradigms: sequencing-based (ST, Visium, Visium HD) and imaging-based (Xenium).
[…]
Xenium: Xenium is an imaging-based spatial-profiling technology offering higher resolution compared to its sequencing counterpart, Visium. Xenium imaging-based paradigm breaks free from the resolution limitations of its sequencing-based counterpart, it directly performs in situ RNA capturing on tissue sections by imaging fluorescent RNA markers derived from padlock probes and rolling circle amplification chemistry, hence offering sub-cellular expression resolution. As of 2024, Xenium cannot perform whole transcriptome measurement and is limited to gene panels of up to 580 genes.
Appendix C HEST
C.1 HEST-1k
We provide a comprehensive description of all publicly available and internal cohorts integrated into HEST-1k.
Table A1: HEST-1k data overview. All included samples include a license that allows sharing and redistributing. National Center for Biotechnology Information: NCBI.
Resource Number of datasets Number of samples Size (GB, raw)
10x Genomics 79 97 225
Mendeley 9 118 181
Spatial-Research 4 139 18
Zenodo 3 17 11
NCBI 42 677 265
Internal 3 28 60
Miscellaneous 2 32 65
Table A2: Datasets gathered from 10x Genomics portal. $n$: number of samples in the cohort.
Collection name Organ Technology $\bm{n}$ Num. genes
Adult Mouse Brain (FFPE) Brain Visium 1 19,465
Adult Mouse Brain Coronal Section (Fresh Frozen) 1 Brain Visium 1 32,285
Adult Mouse Brain Coronal Section (Fresh Frozen) 2 Brain Visium 1 32,285
Adult Mouse Kidney (FFPE) Kidney Visium 1 19,465
Adult Mouse Olfactory Bulb Brain Visium 1 32,285
FFPE Human Breast using the Entire Sample Area Breast Xenium 2 541
FFPE Human Breast with Custom Add-on Panel Breast Xenium 2 541
FFPE Human Breast with Pre-designed Panel Breast Xenium 2 541
FFPE Human Pancreas with Xenium Multimodal Cell Segmentation Pancreas Xenium 1 541
Fresh Frozen Mouse Colon with Xenium Multimodal Cell Segmentation Bowel Xenium 1 541
Fresh Frozen Visium on CytAssist: Human Breast Cancer, Probe-Based Whole Transcriptome Profiling Breast Visium 1 18,085
Fresh Frozen Visium on CytAssist: Mouse Brain, Probe-Based Whole Transcriptome Profiling Brain Visium 1 19,465
Human Bone and Bone Marrow Data with Custom Add-on Panel Bone Xenium 3 541
Human Brain Cancer, 11 mm Capture Area (FFPE) Brain Visium 1 18,085
Human Breast Cancer (Block A Section 1) Breast Visium 1 33,538
Human Breast Cancer (Block A Section 2) Breast Visium 1 33,538
Human Breast Cancer: Ductal Carcinoma In Situ, Invasive Carcinoma (FFPE) Breast Visium 1 17,943
Human Breast Cancer: Targeted, Immunology Panel Breast Visium 1 1,056
Human Breast Cancer: Visium Fresh Frozen, Whole Transcriptome Breast Visium 1 36,601
Human Breast Cancer: Whole Transcriptome Analysis Breast Visium 1 32,285
Human Cerebellum: Targeted, Neuroscience Panel Brain Visium 1 1,186
Human Cerebellum: Whole Transcriptome Analysis Brain Visium 1 1,186
Human Cervical Cancer (FFPE) Cervix Visium 1 17,943
Human Colon Preview Data (Xenium Human Colon Gene Expression Panel) Bowel Xenium 2 541
Human Colorectal Cancer, 11 mm Capture Area (FFPE) Bowel Visium 1 18,085
Human Colorectal Cancer: Targeted, Gene Signature Panel Bowel Visium 1 1,142
Human Colorectal Cancer: Whole Transcriptome Analysis Bowel Visium 1 36,601
Human Glioblastoma: Targeted, Pan-Cancer Panel Brain Visium 1 1,253
Human Glioblastoma: Whole Transcriptome Analysis Brain Visium 1 36,601
Human Heart Heart Visium 1 36,601
Human Heart Data with Xenium Human Multi-Tissue and Cancer Panel Heart Xenium 1 541
Human Intestine Cancer (FPPE) Bowel Visium 1 17,943
Human Kidney Preview Data (Xenium Human Multi-Tissue and Cancer Panel) Kidney Xenium 2 541
Human Kidney, 11 mm Capture Area (FFPE) Kidney Visium 1 18,085
Human Liver Data with Xenium Human Multi-Tissue and Cancer Panel Liver Xenium 2 541
Human Lung Cancer (FFPE) Lung Visium 1 18,085
Human Lung Cancer, 11 mm Capture Area (FFPE) Lung Visium 1 18,085
Human Lymph Node Lymph node Visium 1 36,601
Table A3: Datasets gathered from 10x Genomics portal. Continuation.
Collection name Organ Technology $\bm{n}$ Num. genes
Human Ovarian Cancer (FFPE) Ovary Visium 1 17,943
Human Ovarian Cancer, 11 mm Capture Area (FFPE) Ovary Visium 1 18,085
Human Prostate Cancer, Acinar Cell Carcinoma (FFPE) Prostate Visium 1 17,943
Human Prostate Cancer, Adenocarcinoma with Invasive Carcinoma (FFPE) Prostate Visium 1 17,943
Human Skin Data with Xenium Human Multi-Tissue and Cancer Panel Skin Xenium 2 541
Human Skin Preview Data (Xenium Human Skin Gene Expression Panel with Custom Add-On) Skin Xenium 1 541
Human Skin Preview Data (Xenium Human Skin Gene Expression Panel) Skin Xenium 1 541
Human Tonsil Data with Xenium Human Multi-Tissue and Cancer Panel Lymph node Xenium 2 541
Mouse Bone Data with Custom Add-on Panel Bone Xenium 3 541
Mouse Brain Coronal Section 1 (FFPE) Brain Visium 1 19,465
Mouse Brain Coronal Section 2 (FFPE) Brain Visium 1 19,465
Mouse Brain Section (Coronal) Brain Visium 1 31,053
Mouse Brain Serial Section 1 (Sagittal-Anterior) Brain Visium 1 31,053
Mouse Brain Serial Section 1 (Sagittal-Posterior) Brain Visium 1 31,053
Mouse Brain Serial Section 2 (Sagittal-Anterior) Brain Visium 1 31,053
Mouse Brain Serial Section 2 (Sagittal-Posterior) Brain Visium 1 31,053
Mouse Embryo, 11 mm Capture Area (FFPE) Embryo Visium 1 19,465
Mouse Kidney Section (Coronal) Kidney Visium 1 31,053
Mouse Tissue Microarray in 3x3 Layout with 1 mm Edge to Edge Spacing (FFPE) Lung/Brain Visium 1 19,465
Mouse Tissue Microarray in 3x3 Layout with 2 mm Edge to Edge Spacing (FFPE) Lung/Brain Visium 1 19,465
Mouse Tissue Microarray in 5x5 Layout with 1 mm Edge to Edge Spacing (FFPE) Kidney/Brain Visium 1 19,465
Normal Human Prostate (FFPE) Prostate Visium 1 17943
Pancreatic Cancer with Xenium Human Multi-Tissue and Cancer Panel Pancreas Xenium 1 538
Preservation Method Comparison on CytAssist: FFPE Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preservation Method Comparison on CytAssist: Fixed Frozen Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preservation Method Comparison on CytAssist: Fresh Frozen Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preservation Method Comparison on Visium CytAssist: FFPE Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preservation Method Comparison on Visium CytAssist: Fixed Frozen Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preservation Method Comparison on Visium CytAssist: Fresh Frozen Mouse Brain (Sagittal), 11 mm Capture Area Brain Visium 1 19,465
Preview Data: FFPE Human Lung Cancer with Xenium Multimodal Cell Segmentation Lung Xenium 1 541
Visium CytAssist Gene Expression Libraries of Post-Xenium Human Colon Cancer (FFPE) Bowel Visium 4 18,085
Visium CytAssist Gene Expression Libraries of Post-Xenium Mouse Brain (FF) Brain Visium 4 19,465
Visium CytAssist, Mouse Embryo, 11 mm Capture Area (FFPE) Embryo Visium 1 19,465
Whole Mouse Pup Preview Data (Xenium Mouse Tissue Atlassing Panel) Whole organism Xenium 1 541
Table A4: Datasets gathered from NCBI.
Publication Organ Technology $\bm{n}$ Num. genes
10X Visium Spatial transcriptomics of murine colon at d14 (mucosa healing) in B cell sufficient-deficient mice Parigi et al. (2022) Bowel Visium 2 31,053
10X Visium Spatial transcriptomics of murine colon in steady state and during recovery after DSS colitis Parigi et al. (2022) Bowel Visium 2 31,053
A Spatial Transcriptomic atlas of the human kidney papilla identifies significant immune injury in patients with stone disease Canela et al. (2023) Kidney Visium 7 36,601
A cellular hierarchy in melanoma uncouples growth and metastasis Karras et al. (2022) Skin Visium 3 31,053
A new epithelial cell subpopulation predicts response to surgery, chemotherapy, and immunotherapy in bladder cancer Abdel-Hafiz et al. (2023); Gouin et al. (2021) Bladder Visium 4 33,538
A novel model of binge ethanol exposure reveals enhanced neurodegeneration with advanced age Anton et al. (2024) Brain Visium 4 32,285
A single-cell transcriptomic analysis of endometriosis Fonseca et al. (2023) Uterus Visium 2 36,601
Distinct mesenchymal cell states mediate prostate cancer progression Pakula et al. (2024) Prostate Visium 2 32,589
Epithelial Plasticity and Innate Immune Activation Promote Lung Tissue Remodeling following Respiratory Viral Infection Beppu et al. (2023) Lung Visium 1 32,285
Gene expression within a human choroidal neovascular membrane using spatial transcriptomics Voigt et al. (2023) Eye Visium 5 36,601
Genome-wide Spatial Expression Profiling in Formalin-fixed Tissues Gracia Villacampa et al. (2021) Kidney Visium 14 33,538
High resolution mapping of the tumor microenvironment using integrated single-cell, spatial and in situ analysis Janesick et al. (2023) Breast Xenium 4 541
Identification of TREM1+CD163+ myeloid cells as a deleterious immune subset in HCC Giraud et al. (2022) Liver Visium 2 36,601
Integration of spatial and single cell transcriptomics localizes epithelial-immune cross-talk in kidney injury Ferreira et al. (2021) Kidney Visium 4 33,538
Molecular Atlas of the Adult Mouse Brain Ortiz et al. (2020) Brain Spatial Transcriptomics 75 23,371
Regional differential gene expression analyses of brains from four 24w-old Nf1+- mice Brain Visium 4 32,285
SARS-CoV-2 Niches in Human Placenta Revealed by Spatial Transcriptomics Barrozo et al. (2023) Uterus Visium 16 36,612
Schwann Cells Shape Tumor Cells and Cancer-Associated Fibroblasts in the Pancreatic Ductal Adenocarcinoma Microenvironment Xue et al. (2023) Pancreas Visium 4 20,615
Single Cell and Spatial Analysis of Human Squamous Cell Carcinoma Ji et al. (2020) Skin Spatial Transcriptomics 12 17,138
Single cell profiling of primary and paired metastatic lymph node tumors in breast cancer patients Liu et al. (2022) Lymph node Visium 4 33,931
Table A5: Datasets gathered from NCBI. Continuation.
Publication Organ Technology $\bm{n}$ Num. genes
Single-cell and spatial transcriptomics characterisation of the immunological landscape in the healthy and PSC human liver Andrews et al. (2024) Liver Visium 4 36,601
Single-nucleus Ribonucleic Acid-sequencing and Spatial Transcriptomics Reveal the Cardioprotection of Shexiang Baoxin Pill (MUSKARDIA) in Mice with Myocardial Ischemia-Reperfusion Injury Lin et al. (2023) Heart Visium 2 32,285
Spatial Multimodal Analysis: MALDI-MSI and Spatial Transcriptomics within the same tissue section Vicari et al. (2023) Brain Visium 19 32,285
Spatial RNA sequencing of regenerating mouse hindlimb muscle McKellar et al. (2021) Muscle Visium 3 33,217
Spatial Total RNA-Sequencing of regenerating mouse hindlimb muscle and Type 1-Lang reovirus-infected mouse heart McKellar et al. (2023) Muscle Visium 7 55,414
Spatial localization with Spatial Transcriptomics for an atlas of healthy and injured cell states and niches in the human kidney Lake et al. (2023) Kidney Visium 23 33,538
Spatial sequencing of Foreign body granuloma Krausgruber et al. (2023) None Visium 1 15,524
Spatial transcriptomics landscape of non-communicable inflammatory skin diseases Schäbitz et al. (2022) Skin Visium 59 20,613
Spatial transcriptomics of adenoid cystic carcinoma of the lacrimal gland Moeyersoms et al. (2023) Eye Visium 1 17,943
Spatial transcriptomics of the mouse brain across three age groups Brain Visium 6 32,285
Spatial transcriptomics reveal unnresolved wound repair as potential driver of PFA Ependymoma progression Fu et al. (2023) Brain Visium 14 36,601
Spatiotemporal dynamics of molecular pathology in amyotrophic lateral sclerosis Silas Maniatis et al. (2019) Spinal cord Spatial Transcriptomics 302 12,572
The neurons that restore walking after paralysis Kathe et al. (2022) Spinal cord Visium 16 22,127
Visium spatial transcriptomics analysis of lacrimal gland during chronic inflammation progression Mauduit et al. (2022) Eye Visium 4 32,285
YAP Drives Assembly of a Spatially Colocalized Cellular Triad Required for Heart Renewal Li et al. (2024) Heart Visium 2 32,285
Zika virus co-opts miRNA networks to persist in placental microenvironments detected by spatial transcriptomics Barrozo et al. (2024) Placenta Visium 8 32,298
Mouse model Heptablastoma spatial transcriptomics Pilet et al. (2023) Liver Visium 10 31,053
Table A6: Datasets gathered on Spatial-Research.
Publication Organ Technology $\bm{n}$ Num. genes
A spatiotemporal organ-wide gene expression and cell atlas of the developing human heart Asp et al. (2019) Heart Spatial Transcriptomics 19 39,739
Integrating spatial gene expression and breast tumour morphology via deep learning He et al. (2020) Breast Spatial Transcriptomics 68 16,744
Spatial deconvolution of HER2-positive breast cancer delineates tumor-associated cell type interactions Andersson et al. (2021) Breast Spatial Transcriptomics 36 15,045
Visualization and analysis of gene expression in tissue sections by spatial transcriptomics Ståhl et al. (2016a) Brain Spatial Transcriptomics 16 16,573
Table A7: Datasets gathered on Mendeley.
Publication Organ Technology $\bm{n}$ Num. genes
Ex-ST Andrusivova and Fan (2023) Brain Visium 5 31,053
Genome-wide spatial expression profiling in formalin-fixed tissues Villacampa et al. (2021) Brain Visium 15 31,053
Human ileum, Visium Mirzazadeh et al. (2021) Bowel Visium 4 33,538
Human squamous cell carcinoma Abalo et al. (2021) Skin Visium 4 33,538
Prostate needle biopsies pre- and post-ADT: Count matrices, histological-, and Androgen receptor immunohistochemistry images Marklund (2022) Prostate Spatial Transcriptomics 24 26,437
Spatially resolved clonal copy number alterations in benign and malignant tissue Erickson et al. (2022) Prostate Visium 23 33,538
Spatially resolved transcriptomic profiling of degraded Mirzazadeh et al. (2023) Bowel Visium 35 17,943
spatialRNAseq heart raw suppdata Heart Visium 4 54,848
spatialRNAseq ileum raw suppdata Bowel Visium 4 54,848
Table A8: Datasets gathered on Github and the Human Cell Atlas data explorer.
Publication Organ Technology $\bm{n}$ Num. genes
Transcriptome-scale spatial gene expression in the human dorsolateral prefrontal cortexMaynard et al. (2021) Brain Visium 12 33,538
A spatially resolved atlas of the human lung characterizes a gland-associated immune nicheMadissoon et al. (2023) Lung Visium 20 17,922
Table A9: Internal datasets.
Publication Organ Technology $\bm{n}$ Num. genes
Prostate ST Internal Prostate Visium 4 17,943
Tertiary lymphoid structures generate and propagate anti-tumor antibody-producing plasma cells in renal cell cancer Meylan et al. (2022) Lymph node Visium 24 17,943
Table A10: Datasets gathered on Zenodo.
Publication Organ Technology $\bm{n}$ Num. genes
Charting the Heterogeneity of Colorectal Cancer Consensus Molecular Subtypes using Spatial Transcriptomics: datasets Valdeolivas et al. (2023) Bowel Visium 14 36,601
Demo 10x Visium dataset for STQ zen (2024a) Skin Visium 1 68,886
Spotiphy: generative modeling in single-cell spatial whole transcriptomics zen (2024b) Brain Visium 2 32,285
C.2 HEST-Library
The HEST-Library helps transform unstructured spatial transcriptomics and histology data into a unified format. An overview of the HEST-Library is provided in Suppl. Figure 3. An example of fiducial detection is presented in Suppl. Figure 4.
Figure 3: Overview of HEST-Library functionalities. HEST was designed to transform legacy data scrapped in multiple public repositories, such as NCBI, into unified and actionable objects that can easily be integrated into multimodal computational pipelines.
Figure 4: Fiducial detection and automatic alignment in Visium samples. Corner fiducials on 6.5$\times$6.5mm and 11mm×$\times$11mm Visium slides are automatically detected with a finetuned Yolov8 model, the spots coordinates are then derived if at least 3 of the 4 corner fiducials are detected. This process enables automatically estimating the pixel resolution.
C.3 HEST-Benchmark
Gene selection, Random Forest, and Ridge regression models: We learn a regression model that maps the patch embeddings of each encoder to its corresponding gene expression profile. The Random Forest model is based on 70 trees (sklearn implementation). The Ridge regression uses a fixed $L2$ regularization coefficient $\lambda$ to $100/MC$, where $M$ is the embedding dimension and $C=50$ is the number of targets trained with the Regularized Least-Squares Routine solver (sklearn implementation). We train the regression models to predict a panel of 50 most variable genes tailored to each task. Specifically, for each task, we select the 50 most variable genes across all spots and samples, after excluding the genes that have non-zero counts in less than 10% of the spots.
Benchmark task description: We provide complementary information on each task introduced as part of the HEST-Benchmark.
Table A11: Overview of the HEST-Benchmark. Each task involves predicting the expression levels of the 50 most variable genes from 112$\times$112 $\mu$m H&E-stained image patches centered on each spatial transcriptomics spot. The tasks are formulated as multivariate regression problems. The Oncotree code describes the cancer type diagnosed in samples, e.g., PAAD denotes pancreatic adenocarcinoma. Additional information is provided in the Appendix.
Task ID Oncotree Organ Number of Patients Number of Samples Technology
Task 1 Janesick et al. (2023) IDC Breast 4 4 Xenium
Task 2 PRAD Prostate 2 23 Visium
Task 3 PAAD Pancreas 3 3 Xenium
Task 4 SKCM Skin 2 2 Xenium
Task 5 Valdeolivas et al. (2023) COAD Colon 3 6 Visium
Task 6 Valdeolivas et al. (2023) READ Rectum 2 4 Visium
Task 7 Meylan et al. (2022) ccRCC Kidney 24 24 Visium
Task 8 Giraud et al. (2022) HCC Liver 2 2 Visium
Task 9 LUAD Lung 2 2 Xenium
Task 10 Liu et al. (2022) IDC Axillary lymph nodes 4 4 Visium
Task 1: Expression prediction in invasive ductal carcinoma (breast cancer, IDC). We used all publicly available Xenium samples available on 10x Genomics (“FFPE Human Breast using the Entire Sample Area”, 2 patients) and two samples published in  Janesick et al. (2023) (2 patients). All samples are FFPE sections imaged with the Xenium pipeline v1.
Task 2: Expression prediction in prostate adenocarcinoma (prostate cancer, PRAD). We used all 23 Visium samples (fresh frozen sections) from 2 patients published in Erickson et al. (2022). Both patients were diagnosed with prostatic acinar adenocarcinoma with a (4+3) Gleason score (ISUP group 4).
Task 3: Expression prediction in pancreatic adenocarcinoma (pancreatic cancer, PAAD). We used 3 samples from 3 different patients from 10x Genomics (“FFPE Human Pancreas with Xenium Multimodal Cell Segmentation” and “Pancreatic Cancer with Xenium Human Multi-Tissue and Cancer Panel”). All samples are FFPE sections processed with Xenium pipeline v1.
Task 4: Expression prediction in skin cutaneous melanoma (skin cancer, SKCM). We used 2 samples from 2 different patients from 10x Genomics website (“Human Skin Data with Xenium Human Multi-Tissue and Cancer Panel”). All samples are FFPE sections processed with Xenium pipeline v1.
Task 5: Expression prediction in colon adenocarcinoma (colon cancer, COAD). We used 6 COAD samples from 3 different patients published in Valdeolivas et al. (2023). All samples are fresh frozen sections processed with Visium.
Task 6: Expression prediction in rectal adenocarcinoma (rectum cancer, READ). We used 4 READ samples from different 2 patients published in Valdeolivas et al. (2023). All samples are fresh frozen sections processed with Visium.
Task 7: Expression prediction in clear cell renal cell carcinoma (kidney cancer, ccRCC). We used all 24 ccRCC samples from 24 different patients published in Meylan et al. (2022). All samples are fresh frozen sections processed with Visium.
Task 8: Expression prediction in hepatocellular carcinoma (liver cancer, HCC). We used 2 HCC liver samples from 2 different patients published in Giraud et al. (2022). All samples are fresh frozen sections processed with Visium.
Task 9: Expression prediction in lung adenocarcinoma (lung cancer, LUAD). We used 2 LUAD samples from 2 different patients from 10x genomics (“Preview Data: FFPE Human Lung Cancer with Xenium Multimodal Cell Segmentation”). All samples are fresh frozen sections processed with Xenium pipeline v1.
Task 10: Expression prediction in axillary lymph nodes in IDC patients. We used 4 axillary lymph node samples from 2 IDC patients published in Liu et al. (2022). All samples are fresh frozen sections processed with Visium.
Table A12: State-of-the-art foundation models for histology evaluated on HEST-Benchmark. ResNet-50: 23M parameters, ResNet-18: 11M, ResNet-152: 795M, ViT-B: Base (86M), L: Large (307M), G: Giant (1.13B)
Name Number of Number of Magnification Model Training
slides patches recipe
ResNet50-IN Lu et al. (2021b) N/A 1M N/A ResNet-50 Supervised
KimiaNet Riasatian et al. (2021) 7K 240K 20x ResNet-18 Supervised
Ciga Ciga et al. (2022) UNK UNK 10x, 20x ResNet-50 Supervised
CTransPath Wang et al. (2021) 32k 17M 10x UNK MoCov2
Remedis Azizi et al. (2023) 10 10M 20x ResNet-152 iBOT
Phikon Filiot et al. (2023) 6k 43.3M 20x ViT-B iBOT
PLIP Huang et al. (2023) N/A 800k Many ViT-B CLIP
UNI Chen et al. (2024) 100k 100M 20x ViT-L DINOv2
CONCH Lu et al. (2024) N/A 1.17M Many ViT-B CoCa
GigaPath Xu et al. (2024) 171k 1.3B 20x ViT-G CoCa
We provide a brief description of each patch encoder assessed on the HEST-Benchmark.
ResNet50 (IN) Lu et al. (2020): This model uses a ResNet50 backbone He et al. (2016) trained on ImageNet dataset Deng et al. (2009). Following prior work Lu et al. (2020), the patch embedding corresponds to the deep representation at the penultimate layer before classification.
Ciga Ciga et al. (2022): This model uses a ResNet-50 backbone He et al. (2016) trained on public histology datasets such as TCGA and Cameleyon16 using SimCLR.
CTransPath Wang et al. (2022): This model uses a “Tiny” Swin Transformer backbone Liu et al. (2021) with a window size of 14 (Swin-T/14, 28 million parameters) pretrained on TCGA and PAIP datasets (17 million images) using MoCoV3 Chen et al. (2021).
Remedis Azizi et al. (2023): This model uses a ResNet-152$\times$2 (232 million parameters) initialized with the “Big Transfer”-medium protocol Kolesnikov et al. (2020) on ImageNet-22K and pretrained with SimCLR Chen et al. (2020) on TCGA.
Phikon Filiot et al. (2023): This model uses a Vision Transformer-Base (ViT-B, 86 million parameters) Dosovitskiy et al. (2021) trained on TCGA data using iBOT Zhou et al. (2022).
PLIP Huang et al. (2023): This model uses a ViT-B (86 million parameters) trained on 208 thousand pathology images paired with natural language descriptions extracted from Twitter and trained using CLIP Radford et al. (2021).
UNI Chen et al. (2024): This model uses a ViT-Large (ViT-L, 307 million parameters) Dosovitskiy et al. (2021) trained on 100 million histology images (over 100,000 slides) from proprietary and public data using DINOv2 Oquab et al. (2023).
CONCH Lu et al. (2024): This model uses a ViT-B (86 million parameters) trained on 1.17 million histology image–caption pairs extracted from online educational and research resources using CoCa Yu et al. (2022).
GigaPath Xu et al. (2024): This model uses a ViT-Giant (1.13 billion parameters) trained on 1.3 billion image patches from 171,189 WSIs at 20$\times$ magnification using DINOv2.
Appendix D HEST for multimodal representation learning
CONCH-FT model, a ViT-Base model initialized with CONCH weights, was fine-tuned for 50 epochs using a cosine learning rate scheduler, with a base learning rate of $10^{-4}$ for the image encoder and $10^{-3}$ for the expression encoder. Only the last 3 layers of the model were fine-tuned, with a layer-wise learning decay rate of 0.7. For training with the infoNCE loss, a contrastive temperature of $10^{-2}$ and batch size of 1,024 pairs of patch and transcriptomics were used. A combination of random horizontal/vertical flip and color jittering was employed for image patch augmentation.
The rank of the embedding space, also referred to as smooth rank measure Garrido et al. (2023), measures the quality of the embeddings produced from encoders trained in unsupervised or self-supervised manners. Given the patch embedding matrix $H\in\mathcal{R}^{N\times d}$ and $d<N$, where $N$ is the number of patches and $d$ is the feature dimension, we compute the rank as the entropy of the $d$ L1-normalized singular values of $H$.
Appendix E HEST for discovery
Cells were segmented and classified using CellViT Hörst et al. (2024). To find the gene expression profile of each neoplastic cell, we matched each cell to its corresponding cell index in Xenium by assigning the index for which the distance between the cell centroids was the smallest. After matching all neoplastic cells, only those cells for which the assignment was unique were kept. After this filtering step, an average of 91% of the cells per sample were kept while 9% of the cells were discarded. The discarded cells likely correspond to over-segmentation errors by CellViT.
Figure 5: HEST for biomarker discovery: Analysis of an invasive ductal carcinoma Xenium sample. a. IDC Xenium sample with neoplastic nuclei overlaid in red ($n_{c}$=342,018 detected nuclei). Six randomly selected regions with CellViT segmentation of the neoplastic nuclei. Black scale bar represents 30 $\mu$m. b. Pearson correlation between the major axis length of neoplastic nuclei and the log1p-normalized expression of TPD52. c. Analogous analysis between nuclear area and FLNB expression.
Appendix F Datasheet for HEST-1k
We provide a DataSheet for HEST-1k that summarizes the contributions, analyses, and intended usages presented in the study.
F.1 Motivation for dataset creation
  • Why was the dataset created? HEST-1k was designed with three key applications: (1) multimodal representation learning of histology and transcriptomics, (2) biomarker discovery, and (3) benchmarking of histopathology patch encoder models. Despite many publicly available resources, no existing unified and user-friendly formatting was available to bring ST in the world of deep learning.
  • What (other) tasks could the dataset be used for? Are there obvious tasks for which it should not be used? Users are welcome to propose new creative ways to use the dataset. Users are not allowed to try to retrieve patient information from the existing data. A dedicated section is allocated to discuss ethical considerations and intended usage.
  • Has the dataset been used for any tasks already? If so, where are the results so others can compare (e.g., links to published papers)? All samples that were made public as part of a publication (peer-review or not) are reported in the metadata attached with HEST-1k.
  • Who funded the creation of the dataset? The work was funded by an R35 grant.
F.2 Dataset composition
  • What are the instances? The modalities used in this study are histopathology whole-slide images, gene expression data, and derivatives of these two modalities, such as nuclear segmentation maps.
  • Are relationships between instances made explicit in the data Each whole-slide image maps to a unique gene expression profile in an unequivocal way.
  • What data does each instance consist of? Imaging data consists of Generic TIFF objects stored in pyramidal format, and gene expression data consists of scanpy objects. Derivatives are stored in JSON and Hierarchical Data Format (HDF) files.
  • Is there a label/target associated with instances? If the instances are related to people, are subpopulations identified (e.g., by age, gender, etc.), and what is their distribution? Each sample pair (slide and expression profile) is associated with comprehensive metadata. All metadata information is thoroughly described in the main paper. Age and gender are only reported in a subset of cases.
  • Is everything included or does the data rely on external resources? (e.g., websites, tweets, datasets) If external resources, a) are there guarantees that they will exist, and remain constant, over time; b) is there an official archival version. Are there licenses, fees or rights associated with any of the data? We provide all data as part of HEST-1k release. In addition, a link to the original data is provided in the metadata. Each sample is associated with a license as provided by the original publication, where we ensured that the reported license allowed for distributing and creating derivatives of the data. Should the author of the original publication would like to remove their samples from HEST-1k, we will update the data promptly.
  • Are there recommended data splits or evaluation measures? HEST-1k comes with the HEST-Benchmark, a series of tasks for gene expression prediction from histology images. All patient-stratified splits are specified in the attached comma-separated values (CSV) files.
  • What experiments were initially run on this dataset? Have a summary of those results and, if available, provide the link to a paper with more information here. At the time of submission, all experiments run with HEST-1k are part of this study. The reader can refer to the main text for a thorough description of all experiments (see HEST-Benchmark, HEST for discovery, HEST for multimodal representation learning).
F.3 Data collection process
  • How was the data collected? The data were manually inspected and curated by the authors of the present study.
  • Who was involved in the data collection process? All authors of the present study were involved in the data collection, inspection, and curation. The reader can refer to the original publication to understand how the data were orginally acquired.
  • Over what time-frame was the data collected? Does the collection time-frame match the creation time-frame? The original data comprise publications from 2016 to 2024. As the dataset grows, more recent publications might be included in HEST-1k.
  • Does the dataset contain all possible instances? Or is it, for instance, a sample (not necessarily random) from a larger set of instances? All pairs of gene expression data and whole-slide images of the underlying studies were included and are unique.
  • Is there information missing from the dataset and why? (this does not include intentionally dropped instances; it might include, e.g., redacted text, and withheld documents) Is this data missing because it was unavailable? Original publications include some missing information, such as the alignment file between the slide and the expression profile. We developed computational tools to minimize missing information to reach near-complete metadata.
  • Are there any known errors, sources of noise, or redundancies in the data? All whole-slide images have been manually inspected. The quality from one sample to another varies significantly, for instance, due to poor staining, compression artifact, lower resolution, etc. Gene expression data are inherently noisy. Users can decide to apply post-hoc normalization methods to reduce noise.
F.4 Data preprocessing
  • What preprocessing/cleaning was done? All whole-slide images were converted into a pyramidal TIFF objects with re-estimated pixel resolution. All alignment files have been manually inspected and included if missing. All gene expression data have been transformed into scanpy objects following the same process.
  • Was the “raw” data saved in addition to the preprocessed/cleaned data? (e.g., to support unanticipated future uses) Raw data are downloaded, but not publicly shared. In the case of public samples, users can re-download them using the metadata provided as part of the dataset release.
  • Is the preprocessing software available? Yes, the source code to preprocess HEST-1k is made publicly available as part of the HEST library.
F.5 Dataset distribution
  • How is the dataset distributed? HEST-1k is distributed using internal storage. We have several internal copies to ensure data preservation over the years.
  • When will the dataset be released/first distributed? The dataset will be made public with the camera-ready version of this publication. This study serves as the canonical reference.
  • What license (if any) is it distributed under? Are there any copyrights on the data? The dataset is distributed under the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 Deed).
  • Are there any fees or access/export restrictions? No access/export restrictions unless they violates the terms of the above-mentioned license (CC BY-NC-SA 4.0 Deed).
  • Who is supporting/hosting/maintaining the dataset? How does one contact the owner/curator/manager of the dataset? The dataset is maintained by the authors of the publication.
  • Will the dataset be updated? How often and by whom? How will updates/revisions be documented and communicated (e.g., mailing list, GitHub)? Is there an erratum? The dataset might evolve as additional samples become publicly available. Dataset versioning will be put in place. A website will clearly state what the latest version of the dataset contains and keep track of the changes in each version.
  • If the dataset becomes obsolete how will this be communicated? The corresponding website will state if the dataset becomes obsolete.
  • Is there a repository to link to any/all papers/systems that use this dataset? There is no repository to link papers that use HEST-1k. Users are required to cite HEST-1k if they use it in their own research.
  • If others want to extend/augment/build on this dataset, is there a mechanism for them to do so? If so, is there a process for tracking/assessing the quality of those contributions. What is the process for communicating/distributing these contributions to users? Users are welcome to contact us if they would like to provide additional data that meet the standards of the data contained in HEST-1k. We do not have a dedicated system to communicate these contributions. Newly added data will be tracked in the versioning.
F.6 Legal and ethical considerations
  • If the dataset relates to people (e.g., their attributes) or was generated by people, were they informed about the data collection? (e.g., datasets that collect writing, photos, interactions, transactions, etc.) HEST-1k does not include patient information (such as name, address, etc.).
  • If it relates to other ethically protected subjects, have appropriate obligations been met? (e.g., medical data might include information collected from animals) For animal samples (Mus musculus tissue), we refer to the original publication for an in-depth analysis.
  • If it relates to people, were there any ethical review applications/reviews/approvals? (e.g. Institutional Review Board applications) For human tissue, we refer to the original publication for an in-depth analysis. Internal cohorts were ethically reviewed and collected as part of dedicated IRBs.
  • If it relates to people, could this dataset expose people to harm or legal action? (e.g., financial social or otherwise) What was done to mitigate or reduce the potential for harm? No, patients cannot be linked to the corresponding histology and gene expression profile.
  • If it relates to people, does it unfairly advantage or disadvantage a particular social group? In what ways? How was this mitigated? Most datasets do not include specific demographics. When reported, we include this information in the metadata associated with each sample. To our knowledge, the representation of HEST-1k does not unfairly advantage or disadvantage a particular social group.
  • Does the dataset contain information that might be considered sensitive or confidential? (e.g., personally identifying information) No.
  • Does the dataset contain information that might be considered inappropriate or offensive? No.
Appendix G Author statement
The authors of this paper bear all responsibility in case of violation of rights associated with HEST-1k, HEST-Library, and HEST-Benchmark.
Table A13: HEST-Benchmark results. Ridge regression evaluation on public state-of-the-art patch encoders. Model performance measured with Pearson correlation. We report the mean $\pm$ standard deviation computed over all folds (or patients). Best is bold, second best is underlined.
ResNet50 KimiaNet Ciga CTransPath Remedis Phikon PLIP UNI CONCH Gigapath
IDC 0.446 0.455 0.425 0.499 0.496 0.526 0.450 $\bm{0.544}$ 0.541 $\underline{0.542}$
±$\pm 0.04$$\pm 0.03$$\pm 0.04$$\pm 0.06$$\pm 0.07$$\pm 0.08$$\pm 0.04$$\pm 0.07$$\pm 0.07$$\pm 0.09$
PRAD 0.276 0.323 0.258 $0.290$ 0.264 0.249 0.262 0.257 $\bm{0.353}$ 0.301
±$\pm 0.06$$\pm 0.04$$\pm 0.06$$\pm 0.07$$\pm 0.08$$\pm 0.13$$\pm 0.02$$\pm 0.08$$\pm 0.01$$\pm 0.03$
PAAD 0.343 0.374 0.364 $\underline{0.383}$ 0.288 0.360 0.358 0.377 $\bm{0.424}$ 0.323
±$\pm 0.07$$\pm 0.08$$\pm 0.07$$\pm 0.07$$\pm 0.05$$\pm 0.07$$\pm 0.06$$\pm 0.06$$\pm 0.04$$\pm 0.00$
SKCM $\underline{0.413}$ 0.357 0.381 0.404 0.412 0.369 0.337 0.343 $\bm{0.546}$ 0.190
±$\pm 0.08$$\pm 0.05$$\pm 0.06$$\pm 0.07$$\pm 0.04$$\pm 0.11$$\pm 0.08$$\pm 0.06$$\pm 0.05$$\pm 0.05$
COAD $\underline{0.116}$ 0.103 0.093 0.113 0.075 0.114 0.108 0.110 $\bm{0.135}$ 0.085
±$\pm 0.05$$\pm 0.04$$\pm 0.02$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.05$$\pm 0.06$$\pm 0.06$$\pm 0.03$
READ 0.067 0.058 0.081 0.090 0.077 0.114 0.084 0.124 $\bm{0.154}$ $\underline{0.125}$
±$\pm 0.06$ 0.05 ±$\pm 0.05$$\pm 0.08$$\pm 0.07$$\pm 0.07$$\pm 0.06$$\pm 0.04$$\pm 0.04$$\pm 0.05$
ccRCC 0.210 0.190 0.180 0.214 0.219 $\bm{0.253}$ 0.174 $\underline{0.240}$ 0.221 0.226
±$\pm 0.05$$\pm 0.06$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.05$
HCC $\underline{0.039}$ 0.031 0.027 0.024 0.016 0.038 0.032 0.017 $\bm{0.046}$ 0.008
±$\pm 0.02$$\pm 0.01$$\pm 0.02$$\pm 0.02$$\pm 8.71$$\pm 0.01$$\pm 0.03$$\pm 0.00$$\pm 0.01$$\pm 0.00$
LUAD 0.400 0.419 0.417 0.403 0.311 $\underline{0.422}$ 0.406 0.371 $\bm{0.493}$ 0.297
±$\pm 0.06$$\pm 0.07$$\pm 0.04$$\pm 0.07$$\pm 0.04$$\pm 0.06$$\pm 0.01$$\pm 0.11$$\pm 0.03$$\pm 0.08$
LYM-IDC 0.203 0.227 0.218 0.209 0.167 0.215 0.215 0.224 $\bm{0.247}$ 0.191
±$\pm 0.05$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.04$$\pm 0.05$$\pm 0.03$$\pm 0.04$$\pm 0.02$
Average 0.251 0.253 0.244 0.263 0.232 0.266 0.243 0.260 $\bm{0.316}$ 0.229